Suppose that X1, ... , Xn form a random sample from a distribution for the pdf f(x0) is as follows: 1 f(x|0) = - 1x-01 for 0 < x < 0 2 Also suppose that the value of 0 is unknown (-<< 0). Find the MLE of 0.


1
Expert's answer
2021-04-15T16:56:36-0400

The pdf of each observation has the following form:

"f(x|θ) = (\dfrac{1}{θ}, for 0 ≤ x ≤ θ"

      "(0, otherwise"


Therefore, the likelihood function has the form


"L(θ) = ( \dfrac{1}{θ^n} , \text{ for } 0 ≤ x_i ≤ θ (i = 1, · · · , n)"

      "( 0 , otherwise"


It can be seen that the MLE of θ must be a value of θ for which "θ ≥ x_i, \text{ for } i = 1, · · · , n" and

which maximizes "\dfrac{1}{θ^n}" among all such values. Since "\dfrac{1}{θ^n}" is a decreasing function of θ, the estimate will be the smallest possible value of θ such that "θ ≥ x_i \text{ for } i = 1, · · · , n" . This value

is "θ = \text{ max }(x_1, · · · , x_n)" , it follows that the MLE of θ is "\hat{θ} = max(X_1, · · · , X_n)"